OpenStudy (anonymous):

\[(-3x ^{-3}y)^{3}\]

OpenStudy (anonymous):

\[=(-1)^327(x^{-3})^{3}y^3 = -27x^{-9}y^3 = -\frac{27y^3}{x^9}\]